A car of 1500kg tied to the rope of 7m whose other end is located at the centre. If the car is moving at the rate of 5m/s . Determine the centripetal force acting on it


1
Expert's answer
2021-05-20T10:08:49-0400

"F=m\frac{v^2}{r}=1500\cdot\frac{5^2}{7}=5357\ (N)" . Answer